# A Comprehensive Evaluation and Practice of System Penetration Testing

### v1

@@ 4.4.2 SQL Injection Attacks
Most common web applications store user data in databases. SQL injection is a widely used method for testing database security. Its core principle involves entering maliciously crafted SQL code into user input fields to alter query logic, thereby bypassing authentication mechanisms and executing malicious operations on the database. This exploit arises when applications fail to adequately filter user input, allowing attackers to use techniques such as single-quote closing, union queries, and Boolean blind injection to probe database structures and extract critical data. The following experiment provides an intuitive understanding of SQL injection.
@@ 4.4.3 Cross-Site Scripting (XSS) Attacks
Attackers inject malicious code into web pages. When users access pages containing these code fragments, the malicious code executes. The main types are as follows.
@@ 5.1 Paris Olympics Cyberattack Incident
During the 2024 Paris Olympics, French authorities reported over 140 cyberattacks, primarily targeting event-related organizations. Most victims experienced system outages, while a minority suffered server paralysis from DDoS attacks. Other incidents involved system intrusions and data theft. The 2024 Paris Olympics Infrastructure Attack Report released by BforeAI identified approximately 166 domains exploited by attackers for criminal activities. Primary tactics included DNS abuse attacks such as keyword stuffing, impersonating brands for fraud, and phishing emails disguised as Olympic Committee notifications to lure staff into clicking malicious links.
@@ 5.2 AT&T Data Breach Incident
In July 2024, data hosted by AT&T on a third-party cloud service provider was compromised, affecting approximately 110 million customers. To prevent the stolen data from being publicly disclosed, AT&T ultimately paid hackers a ransom of approximately \$370,000. The attack succeeded because the cloud provider’s API interfaces were improperly configured, allowing attackers to bypass authentication and bulk-export user data. The leaked data also suggested that internal employee accounts might have been compromised through phishing or theft. The most critical factor was that sensitive fields, such as user Social Security numbers and password hashes, were stored in plaintext or weakly encrypted formats, and log monitoring failed. Attackers exported data for months without triggering anomaly detection systems. After the incident, the company terminated its partnership with the third-party provider and migrated the data back to its internal platform. The company also forced all users to reset their passwords and replaced the original MD5 hash algorithm with the bcrypt encryption algorithm.
@@ 5.3 Ivanti VPN Zero-Day Exploit Incident
Ivanti Connect Secure VPN is an enterprise-grade remote access solution offering secure remote access and multi-factor authentication capabilities. In January 2024, two zero-day vulnerabilities were disclosed: CVE-2024-21887 (a command injection vulnerability enabling remote code execution, CVSS score 9.1) and CVE-2024-21893 (attacks via forged server-side requests, CVSS score 8.8). It took three weeks for Ivanti to release patches, during which attacks caused significant damage to government agencies, healthcare, and financial sectors. Excessively long patch release intervals allowed attackers to establish persistent access channels. Moreover, most victims did not enable audit logs on their VPN devices, making it impossible to accurately trace the attack path.

### v2

@@ 4.4.2 SQL Injection Attacks
Most common web applications store user data in databases. SQL injection is a widely used method for testing database security. Its core principle involves entering maliciously crafted SQL code into user input fields to alter query logic, thereby bypassing authentication mechanisms and executing malicious operations on the database [6]. This exploit arises when applications fail to adequately filter user input, allowing attackers to use techniques such as single-quote closing, union queries, and Boolean blind injection to probe database structures and extract critical data. The following experiment provides an intuitive understanding of SQL injection.
@@ 4.4.3 Cross-Site Scripting (XSS) Attacks
Attackers inject malicious code into web pages. When users access pages containing these code fragments, the malicious code executes. The main types are as follows [40].
@@ 5.1 Paris Olympics Cyberattack Incident
During the 2024 Paris Olympics, French authorities reported over 140 cyberattacks, primarily targeting event-related organizations. Most victims experienced system outages, while a minority suffered server paralysis from DDoS attacks. Other incidents involved system intrusions and data theft. The 2024 Paris Olympics Infrastructure Attack Report released by BforeAI identified approximately 166 domains exploited by attackers for criminal activities. Primary tactics included DNS abuse attacks such as keyword stuffing, impersonating brands for fraud, and phishing emails disguised as Olympic Committee notifications to lure staff into clicking malicious links [57].
@@ 5.2 AT&T Data Breach Incident
In July 2024, data hosted by AT&T on a third-party cloud service provider was compromised, affecting approximately 110 million customers [10]. To prevent the stolen data from being publicly disclosed, AT&T ultimately paid hackers a ransom of approximately \$370,000. The attack succeeded because the cloud provider’s API interfaces were improperly configured, allowing attackers to bypass authentication and bulk-export user data. The leaked data also suggested that internal employee accounts might have been compromised through phishing or theft. The most critical factor was that sensitive fields, such as user Social Security numbers and password hashes, were stored in plaintext or weakly encrypted formats, and log monitoring failed. Attackers exported data for months without triggering anomaly detection systems. After the incident, the company terminated its partnership with the third-party provider and migrated the data back to its internal platform. The company also forced all users to reset their passwords and replaced the original MD5 hash algorithm with the bcrypt encryption algorithm [23].
@@ 5.3 Ivanti VPN Zero-Day Exploit Incident
Ivanti Connect Secure VPN is an enterprise-grade remote access solution offering secure remote access and multi-factor authentication capabilities [5]. In January 2024, two zero-day vulnerabilities were disclosed: CVE-2024-21887 (a command injection vulnerability enabling remote code execution, CVSS score 9.1) and CVE-2024-21893 (attacks via forged server-side requests, CVSS score 8.8). It took three weeks for Ivanti to release patches, during which attacks caused significant damage to government agencies, healthcare, and financial sectors. Excessively long patch release intervals allowed attackers to establish persistent access channels. Moreover, most victims did not enable audit logs on their VPN devices, making it impossible to accurately trace the attack path.